Sucrose (table sugar) has a molar mass of 342.30 g/mole.

342.30 g/mole. How many sucrose molecules are in a 8.80-pound

8.80-pound bag of sugar? (1 pound=453.6 grams.)



1
Expert's answer
2022-03-31T00:08:05-0400

1pound = 453.6g

8.8× 453.6= 3991.68g

Moles= mass/ RMM

= 3991.68/342.3= 11.66molecules